OpenStudy (emilyjones284):

MEDAL AND FAN! (: If f(x) = 2x – 7, what is the equation for f^–1(x)?

jimthompson5910 (jim_thompson5910):

Were you able to get anywhere with this one?

OpenStudy (emilyjones284):

nope

jimthompson5910 (jim_thompson5910):

alright I'll show you how to do this so you can apply it to others

OpenStudy (emilyjones284):

okay thanks

jimthompson5910 (jim_thompson5910):

Step 1) Replace f(x) with y Step 2) Swap x and y Step 3) Solve for y \[\large f(x) = 2x – 7\] \[\large y = 2x – 7\] \[\large x = 2y – 7\] \[\large x+7 = 2y\] \[\large 2y = x+7\] \[\large y = \frac{x+7}{2}\] \[\large f^{-1}(x) = \frac{x+7}{2}\]

jimthompson5910 (jim_thompson5910):

hopefully that helps and makes sense (and that you can use this on future similar problems)

OpenStudy (emilyjones284):

That looks easy enough.. thanks because i have another problem just like that (:

jimthompson5910 (jim_thompson5910):

ok great, hopefully you will know how to solve it